# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

### v2

@@ Abstract
Generative Artificial Intelligence (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation. We define the GenAI Generation as a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Notably, readiness appears increasingly tied to exposure to GenAI through one’s coursework. Students with greater curricular exposure to GenAI tend to feel more prepared, while those without it more often express vulnerability and uncertainty, highlighting a new and growing divide in readiness that goes beyond traditional disciplinary boundaries. Evaluation of more than 250 responses, with over 40% providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts. The challenge ahead involves implementing associated recommendations for educational institutions, moving beyond the baseline of access toward more informed guidance on the use of these tools, while preserving critical thinking, ethical reasoning, and adaptive learning.
Keywords: Generative Artificial Intelligence; GenAI Generation; Student Perceptions; Higher Education; Career Readiness; Cognitive Development; Adaptability; Problem-Solving; Critical Thinking; Ethics in AI
1 Introduction
The significant advancement of Artificial Intelligence (AI), particularly Generative AI (GenAI), is reshaping the educational landscape. GenAI is a type of AI trained on large datasets to generate new content such as text, images, audio, or video in response to user prompts, setting it apart from predictive models by its ability to create [22, 28, 20, 21]. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI continues to evolve exponentially and integrate within academic contexts, there is a pressing need to understand how it is shaping the developmental experiences of students in their formative academic years. Unlike previous waves of technology’s impact on education, GenAI’s pace and scale introduce a layer of uncertainty around its long-term effects on cognition, critical thinking, and engagement. By honing in on student perspectives, this research provides early insight into how GenAI is influencing not only perceived career readiness but also the cognitive processes underpinning learning and decision-making.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, it provides actionable insights into how the GenAI Generation—students most directly impacted by the rise of GenAI—understand its role in shaping both their academic experiences and future trajectories.
The GenAI Generation
@@ 1 Introduction
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [11, 5, 16, 4]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [4, 5]. This study introduces the GenAI Generation. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
  • Developmental Context: Actively undergoing critical cognitive and social growth while navigating the GenAI Revolution.
  • Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies.
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a highly dynamic environment.
  • Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
Figure 1: Illustrates the relationships between key characteristics, emphasizing the ongoing feedback loop that drives development and adaptation amidst uncertainty.
@@
Just as Google transformed how students accessed and retained information, leading to phenomena dubbed the “Google Effect” [26], and the rise of digital technologies sparked concerns about “digital dementia” [27], the GenAI Revolution represents another critical shift in the educational landscape. These past transformations demonstrate how new technologies can reshape cognitive processes, social behaviors, and academic growth, sometimes with unintended, long-term consequences [25, 6, 18]. The GenAI Generation, defined by its unique characteristics, is similarly at the forefront of this era’s challenges and opportunities. Evaluating this cohort is essential to understand how these tools influence learning and development during critical formative years, ensuring that we can build a foundation of guidance, policies, and research to support them and future generations effectively.
@@
For the purpose of this study, students’ awareness of GenAI means their familiarity with tools such as ChatGPT and their understanding of how these tools can be leveraged in educational and professional contexts. Research has revealed varying levels of awareness globally. One study found widespread awareness among students across 76 countries, with significant usage for academic tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity persist [29]. Another study highlighted differences in engagement with GenAI based on demographic factors, underscoring the need for targeted educational support [10]. Research also shows that while GenAI is valued for idea generation and brainstorming, there are reservations regarding its use in completing assignments, emphasizing the need for ethical policies [3].
The transformative potential of GenAI in education has been widely discussed, particularly in its applications for personalized learning and adaptive tutoring systems, which support higher-order cognitive tasks such as problem-solving and critical analysis [8]. However, challenges like data privacy and instructor resistance remain significant barriers to effective integration. Other studies have explored how personalized learning and language support tools can aid students while also warning of potential over-reliance and its impact on critical thinking [17].
Current research on student awareness of GenAI’s impacts reveals perspectives shaped by demographic and geographic factors. These findings highlight the importance of adopting diverse approaches to educate students about GenAI while providing guidance on data privacy and ethical considerations.
@@
For the purpose of this study, student preparedness refers to their confidence in navigating a dynamic, AI-driven job market and their perception of how well their education equips them for this future. Research highlights gaps in AI-specific training within higher education curricula, which leave students feeling under-prepared for the evolving workforce [1, 7]. Ethical and accuracy concerns are also noted as barriers to effective integration, further underscoring the importance of incorporating AI literacy and practical applications into academic programs to enhance student readiness [7]. Additionally, project-based and interdisciplinary learning approaches are recommended to prepare students to critically engage with AI tools and align their academic training with workforce demands [9].
Other studies propose professional development programs for educators to ensure that GenAI integration aligns with institutional goals and emphasizes critical digital literacy, enabling both students and faculty to navigate the complexities of GenAI in professional environments [23]. Personalized learning supported by GenAI is also identified as a potential way to improve academic and employability outcomes, though concerns about ethical use and critical thinking remain [12].
Students’ preparedness highlights the interdisciplinary importance of addressing GenAI in the classroom, ensuring clarity and intentionality in these discussions. Although this research does not examine faculty perspectives, it is critical to recognize that educators also require support and guidance to effectively integrate GenAI discussions and tools into their curricula.
@@
For the purpose of this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact. Current research highlights ethical issues such as data privacy and bias, alongside practical challenges like overreliance and the risk of misinformation. Studies have noted strong reservations about using GenAI for completing assignments without disclosure as well as concerns about the uncritical acceptance of AI outputs [3, 2]. Additionally, research underscores how the misuse of GenAI could undermine academic integrity and originality, emphasizing the need for critical evaluation skills and clear institutional guidelines for responsible use [29, 2]. Balancing the benefits of automation with safeguards against biases and ethical risks remains a pressing concern.
Some researchers critique traditional assessment models, advocating for redesigned evaluations that prioritize originality and critical engagement [19]. Others warn that over-reliance on GenAI could diminish critical thinking and intellectual engagement, urging for its balanced integration into educational practices [17]. While GenAI tools have the potential to enhance advanced cognitive processes, concerns persist regarding their impact on foundational learning skills such as critical thinking and engagement [3, 29, 2, 19, 17].
This contrast highlights the urgent need for clear guidance on student use of GenAI across all educational levels. Strategies must be developed to ensure that these tools are employed ethically while fostering constructive critical engagement and safeguarding cognitive development.
@@
This study offers timely insights into how students in their formative educational years are experiencing the rapid rise of Generative AI. It introduces the term “GenAI Generation” to describe students navigating both academic development and technological transformation without consistent institutional guidance or established norms.
Drawing from a combination of structured survey data and rich qualitative responses, the study reveals meaningful patterns in students’ awareness, preparedness, and concern regarding GenAI. Across majors, students expressed a desire for more intentional support, not just access to GenAI tools, but structured opportunities to learn how to use them responsibly. While formal curriculum integration remains inconsistent, students who encounter GenAI through coursework or self-directed exploration are more likely to articulate confidence and optimism about its role in their future.
From enthusiasm to anxiety, the findings illuminate the psychosocial and pedagogical dimensions of GenAI’s presence in higher education. Students are not only engaging with a new technology, they are forming perceptions of their own readiness, trust in institutions, and ethical responsibility in real time. These perspectives carry clear implications for curriculum design, faculty preparation, and equitable access to digital literacy.
This work contributes to an emerging understanding of what students need to thrive in a GenAI-influenced world: not only technical fluency, but critical engagement, ethical grounding, and institutional clarity. The study underscores the importance of proactive educational responses that extend beyond STEM fields to support students across all disciplines in this pivotal moment.
2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. This study received Institutional Review Board (IRB) approval from the [Institution Name Redacted for Review], and all participants provided informed consent prior to beginning the survey. Designed for completion in under five minutes [13], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [24]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [14], this was an intentional design choice to encourage engagement and reduce bias [15]. The survey was administered across two [Institution Name Redacted for Review] student cohorts, with full question text available in the supplemental materials.
The first cohort, Fall 2023 Computer Science Graduate students, survey was distributed to Online Computer Science Master’s (OMSCS) students, approximately 12,000 members [CITE REDACTED FOR REVIEW], and campus Computer Science Master’s (MSCS), approximately 14,500 members [CITE REDACTED FOR REVIEW], via listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1.
The second cohort, Fall 2023 Undergraduate students, survey was conducted with support from the [Institution Name Redacted for Review] Office of Academic Effectiveness (OAE), which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The demographic breakdown of the sample can be found in the supplemental materials. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Analysis Approach
Descriptive statistics and frequency distributions were used to compare response patterns across student subgroups, while qualitative evaluation of open-ended responses enabled a systematic exploration of emerging themes and perceptions.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Where the first cohort of graduate students comprised soley of computer science majors, the undergraduate cohort had many degrees. Rather than broadly distinguishing between STEM and non-STEM, degrees were categorized into Computer Science, Engineering, and Other, based on disciplinary focus and methodological orientation. Table 1 provides a Major-Category mapping for reference. This grouping structure reflects both practical sample sizes (Other, 28 Computer Science, and 60 Engineering) and a meaningful distinction in how students may engage with GenAI in academic contexts. This also enabled a comparison between computer science graduate and undergraduate perspectives.
Grouping a diverse range of degrees—such as Biology, Business Administration, Literature, and Public Policy—under the single category “Other” introduces a potential limitation, as it may obscure important disciplinary differences in students’ exposure to or perspectives on GenAI. While specific exposure to AI tools and curricula was not measured directly, these categories serve as a reasonable proxy for varying degrees of disciplinary integration with GenAI-related content, allowing exploratory comparisons across educational domains.
Table 1: Undergraduate Sample and Response Rates by Majors and Category
Major* Category Sample Freq. Response Freq. Sample % Response %
Applied Physics Other 12 2 0.2 1.7
Biology Other 162 3 2.9 2.6
Business Administration Other 440 3 7.8 2.6
Computational Media Other 98 1 1.7 0.9
Earth & Atmospheric Sciences Other 13 1 0.2 0.9
Econ & Int’l Affairs Other 15 1 0.3 0.9
Economics Other 49 2 0.9 1.7
History, Technology, & Society Other 18 2 0.3 1.7
Industrial Design Other 88 2 1.6 1.7
Int’l Affairs & Mod Lang Other 18 2 0.3 1.7
Lit., Media, & Communication Other 59 2 1.0 1.7
Mathematics Other 77 3 1.4 2.6
Neuroscience Other 146 3 2.6 2.6
Physics Other 68 5 1.2 4.3
Psychology Other 66 5 1.2 4.3
Public Policy Other 37 1 0.7 0.9
Aerospace Engineering Engineer 349 8 6.2 7.0
Biomedical Engineering Engineer 399 6 7.1 5.2
Chemical and Biomolecular Eng Engineer 187 5 3.3 4.3
Computer Engineering Engineer 187 2 3.3 1.7
Electrical Engineering Engineer 187 3 3.3 2.6
Environmental Engineering Engineer 58 3 1.0 2.6
Industrial Engineering Engineer 344 5 6.1 4.3
Materials Science & Engr Engineer 67 3 1.2 2.6
Mechanical Engineering Engineer 582 12 10.3 10.4
Nuclear & Radiological Engr Engineer 32 2 0.6 1.7
Computer Science Computer Science 1419 28 25.1 24.3
Applied Lang/Intercultural St No Response 4 0 0.1 0.0
Architecture No Response 76 0 1.3 0.0
Atmospheric & Oceanic Sciences No Response 3 0 0.1 0.0
Biochemistry No Response 88 0 1.6 0.0
Building Construction No Response 13 0 0.2 0.0
Chemistry No Response 40 0 0.7 0.0
Environmental Science No Response 10 0 0.2 0.0
Global Econ/Mod Lang No Response 5 0 0.1 0.0
International Affairs No Response 22 0 0.4 0.0
Music Technology No Response 23 0 0.4 0.0
Solid Earth/Planetary Sciences No Response 2 0 0.0 0.0
Special/Non-Degree No Response 64 0 1.1 0.0
Civil Engineering No Response 128 0 2.3 0.0
*For reference; sample was not balanced on this attribute. Sample breakdowns by additional attributes can be found within the supplemental material.
For the undergraduate cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the graduate cohort and 115 for the undergraduate cohort. Table 2 shows this process
Table 2: Table showing data cleaning process
Survey Population Total Started Total Incomplete Total Dropped Reason for Drop Remaining Completed Total
Computer Science Graduates 2023 26,000** 122 19 1 Response was “prefer not to answer” for degree 102
Undergraduates 2023 5,655 143 26 2 Did not have “undergraduate” for degree type 115
** Approximate total population estimate for Computer Science Graduate 2023 cohort.
Qualitative Free-Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
3 Results
The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines. The graduate cohort response count was 102 and comprised solely of computer science majors. The full breakdown between sample to response counts can be seen in Table 1.
@@ 3 Results
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, seen in Figure 3, with undergraduate computer science students slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
@@
Concern: The majority view GenAI’s career impact positively, seen in Figure 7, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with excitement and optimism about positive career impacts, concern was even greater across all undergraduate degree categories. While less, the computer science graduate cohort still showed concern with 65% concerned with GenAI’s.
@@
Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions. These qualitative insights deepen understanding of the perspectives behind the patterns and are summarized below. 40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt. Participation rates for the free response can be seen in Table 3, showing non-Computer Science students having the highest participation. The Fall undergraduate free responses were analyzed as a whole.
Table 3: Optional Free Response Participation by Question & Degree Category
Category Institutional Recommendation Total Additional Thoughts Total Survey Response Total
Undergraduate Other 17 11 38
Undergraduate Engineer 24 11 49
Undergraduate Computer Science 6 5 29
Graduate Computer Science 52 27 102
@@
The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use GenAI tools (Figure 8. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses. There were fewer who responded to the optional ”additional thoughts” question, those who did showed clear emotional tone(Figure 9). Over half of undergraduates expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes(Figure 10). The computer science graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
@@
4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers. Although surveys were distributed uniformly to ensure equal access across cohorts, the potential for selection bias and low response frequencies remains a limitation. Response counts were modest compared to sample size, and some subgroups—particularly within non-CS or less tech-aligned majors—were underrepresented. As a technology-focused institution, [Institution Name Redacted for Review]’s student population may also have greater exposure to emerging technologies, including GenAI, than students at other institutions. Even with these constraints, the findings offer important early insight into how GenAI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.

### v3

@@ Abstract
Generative Artificial Intelligence (Gen AI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, Gen AI has heralded a unique era and given rise to the Gen AI Generation, a cohort of individuals whose development has been increasingly shaped by the opportunities and challenges Gen AI presents during its widespread adoption within society. This study examines higher education students’ perceptions of Gen AI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. ”Readiness” appears increasingly tied to exposure to Gen AI through one’s coursework. Students with greater curricular exposure to Gen AI tend to feel more prepared, while those without more often express vulnerability and uncertainty, highlighting a new and growing divide that goes beyond traditional disciplinary boundaries. Evaluation of more than 250 responses, with over 40 percent providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for Gen AI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of Gen AI for future career impacts.
I Introduction
@@
The Gen AI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [12, 5, 17, 4]. This study introduces the Gen AI Generation. The Gen AI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the experimental, unregulated rise of Gen AI. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
Developmental Context: Actively undergoing critical cognitive and social growth while navigating the Gen AI Revolution. Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies. High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with Gen AI in a highly dynamic environment. Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
@@
Just as Google transformed how students accessed and retained information, dubbed the “Google Effect” [22], and the rise of digital technologies sparked concerns about “digital dementia” [23], the Gen AI Revolution represents another critical shift. Evaluating the Gen AI Generation is essential to understand how these tools influence development, ensuring we can build a foundation of guidance and informed policies.
@@
For this study, students’ awareness of GenAI refers to their familiarity with Gen AI tools and their understanding of how such tools can support academic and professional work. Prior research shows high global awareness and frequent use for tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity remain [24]. Engagement with GenAI also varies across demographic groups, indicating a need for targeted educational support [11]. While students value GenAI for idea generation and brainstorming, many remain hesitant to use it for completing assignments, underscoring the need for clear ethical guidance [3].
Work on GenAI’s educational potential highlights its promise for personalized learning and adaptive tutoring that foster higher-order cognitive skills [7]. Additional studies note that while personalized support tools can aid learning, over-reliance may hinder critical thinking, reinforcing the need for balanced instructional approaches [18]. Current research shows student awareness is shaped by geo-demographic factors, suggesting that diverse and ethically grounded strategies are needed to prepare students for GenAI’s role in education.
@@
For this study, student preparedness refers to their confidence in navigating an AI-driven job market and their sense of how well their education equips them for that future. Research identifies persistent gaps in AI-focused training within higher education, leaving many students feeling under-prepared, especially amid ethical and accuracy concerns that hinder wider integration [1, 6]. Scholars recommend embedding AI literacy, practical applications, and interdisciplinary, project-based learning to better align academic preparation with workforce demands [8].
Faculty readiness is also critical: professional development programs can help ensure GenAI adoption supports institutional goals and strengthens critical digital literacy for both educators and students [20]. Personalized learning supported by GenAI may further enhance academic and employability outcomes, though concerns about ethics and impacts on critical thinking remain [13]. Overall, student preparedness underscores the need for intentional, interdisciplinary engagement with GenAI in the classroom, alongside adequate support for educators integrating these tools.
@@
For this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact. Research highlights ethical risks, such as data privacy, bias, and misinformation, alongside worries about overreliance and uncritical acceptance of AI outputs [3, 2]. Students express particular reservations using GenAI for assignments without disclosure, and studies warn that misuse may undermine academic integrity and originality, underscoring the need for critical evaluation skills and clear institutional guidelines [24, 2] [19]. Related work warns that over-reliance on GenAI could weaken critical thinking and foundational learning skills, calling for balanced and intentional integration into educational practices [18]. While GenAI can support advanced cognitive processes, persistent concerns about ethics, bias, and cognitive impacts highlight the need for coherent guidance that promotes responsible use while safeguarding student development [3, 24, 2, 19, 18].
@@
This study provides insight into how higher education students are experiencing the rapid rise of Gen AI, introducing the term “Gen AI Generation”. Drawing on structured survey data and qualitative responses, the findings reveal patterns in students’ awareness, preparedness, and concern, showing a clear desire for intentional support.
II Methods
The survey included eight main questions, three demographic items, and two open-ended prompts. The study received Georgia Tech (GT) IRB exemption, and all participants provided informed consent. Designed for completion in under five minutes [14], the survey opened with required disclosures and a screening item, progressing from general to more sensitive topics [21]. Items were refined through lab feedback and internal validation. Attitudes were measured using a four-point forced-choice Likert-type scale, intentionally omitting a neutral option to encourage engagement and reduce satisficing [15, 16]. Survey administration and full question wording are provided in the supplemental materials.
Two GT cohorts were sampled. The Fall 2023 Computer Science Graduate cohort included online ($\approx$12,000 students[10]) and on-campus students ($\approx$14,500[9]), recruited via listservs from Sept. 28–Nov. 1, 2023. The Fall 2023 Undergraduate cohort was recruited through the Office of Academic Effectiveness using stratified random sampling of 5,655 undergraduates balanced by gender, race, citizenship, and class standing (2% margin of error). The survey ran Nov. 2–23, 2023 with reminder emails on Nov. 2, 15, and 22. Demographic breakdowns appear in the supplemental materials. Analysis included descriptive statistics and frequency distributions to compare subgroup patterns, supplemented by qualitative analysis of open-ended responses. For qualitative data, preprocessing involved tokenization, lowercasing, removal of punctuation and stop words, and extraction of unigrams through quadgrams with low-frequency terms excluded. Grammatical parsing of noun and verb phrases supported inductive coding to identify recurring themes.
III Results
The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines. The graduate cohort response count was 102 and comprised solely of computer science majors. The full breakdown between sample to response counts can be seen in the supplemental materials.
@@ III Results
Awareness: A small percentage of respondents in all groups believe Gen AI is a short-lived trend, seen in Figure 2, with undergraduate computer science students slightly more likely to agree at 21%. Most respondents view Gen AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively. When asked about the limitations of Gen AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%. Gen AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
@@
Concern: The majority view Gen AI’s career impact positively, seen in Figure 6, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively. Excitement about Gen AI benefits was high across all groups, peaking among Engineering at 77% and Computer Science Grads at 80%. Other respondents notably lower at 67%. Even with excitement and optimism about positive career impacts, concern was even greater across all undergraduate degree categories. While less, the computer science graduate cohort still showed concern with 65% concerned with Gen AI’s.
@@
Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions. These qualitative insights deepen understanding of the perspectives behind the patterns. 40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt.
@@
The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use Gen AI tools (Figure 7). About 33% of respondents also suggested integrating Gen AI into existing curricula or adding new courses. Ethics and career implications of Gen AI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
There were fewer who responded to the optional additional thoughts question, those who did showed clear emotional tone(Figure 8). Over half of undergraduate free responses expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to specifically express positivity. Additional thoughts covered a broader range of themes(Figure 9). The computer science graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
@@
IV Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Although surveys were distributed uniformly, the potential for selection bias and low response frequencies remains a limitation. Response counts were modest compared to sample size, and some subgroups, particularly within non-STEM majors, were underrepresented. As a technology-focused institution, Georgia Tech’s student population may also have greater exposure to emerging technologies, including Gen AI, than students at other institutions. Even with these constraints, the findings offer important early insight into how Gen AI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.